OpenStudy (anonymous):

the region enclosed by the graphs of y=e^(x/2), y=1, and x=ln3 is revolved about the x-axis. Find the volume of the solid generated??????? I NEED CALCULUS HELP PLEASE

OpenStudy (amistre64):

yay!! an easy one

OpenStudy (anonymous):

easy? lol i would love your help then :)

OpenStudy (amistre64):

lets see where these curves meet; and what our boundaries are

OpenStudy (amistre64):

e^(x/2) = 1 when x=0 right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

x = ln(3) is just a constant; we can find it on a calculator

OpenStudy (anonymous):

yea 1.098

OpenStudy (amistre64):

right; 1.09861...

OpenStudy (amistre64):

let me draw a pic of my interpretation of the bonds then

OpenStudy (anonymous):

i think they intersect at 0 and ln3 but i dont know how to integrate it

OpenStudy (amistre64):

like this?

OpenStudy (amistre64):

theres only two things to integrate here; from 0 to ln(3) right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

lets find the volume of the top curve; then we will cut out the section below y=1ok

OpenStudy (anonymous):

yes that would work but im suppose to do something with (pie)

OpenStudy (amistre64):

pi {S} [e^(1/2(x))]^2 dx ; [0,ln(3)]

OpenStudy (amistre64):

what is e^(x/2)^2 =? e^x right?

OpenStudy (anonymous):

yes because they would cancel

OpenStudy (amistre64):

the key here is to realize that we are adding up all the areas of circles that have been sliced; the area of a circle = pi r^r it just so happens the r = f(x)

OpenStudy (amistre64):

r^2 that is lol

OpenStudy (amistre64):

so; [f(x)]^2 = e^(x/2)^2 = e^x pi {S} e^x dx ; [0,ln(3)] whats the integral of e^x?

OpenStudy (anonymous):

the antiderivetive? its justs e^x

OpenStudy (amistre64):

e^x is its own derivative; so its its own integral :) yes

OpenStudy (amistre64):

pi [ e^(ln(3)) e^(0) ] = ?

OpenStudy (amistre64):

typoed that

OpenStudy (amistre64):

pi [ e^(ln(3)) - e^(0) ] = ?

OpenStudy (anonymous):

an example my teacher did shows it being {S} (pi (e^x)-1) and then taking the antiderivetive of that

OpenStudy (amistre64):

lets stay on one track here; well jump the track when we get to the end

OpenStudy (anonymous):

ok i understand that i just dont understand where the minus 1 came from?

OpenStudy (amistre64):

right now we are making a solid piece; we will then cut out the senter

OpenStudy (amistre64):

pi [e^(ln3) - e^0] = ??

OpenStudy (anonymous):

o so we r taking the whole volume minus the part we dont need which is the y=1

OpenStudy (amistre64):

exaclty :) we can do it in parts or all togther; makes no difference

OpenStudy (anonymous):

gotcha and i got 2 for that answer

OpenStudy (amistre64):

2 or 2pi?

OpenStudy (anonymous):

2 pi

OpenStudy (amistre64):

good :) now let integrate y = 1 from 0 to ln(3); or do you know a shortcut for this one?

OpenStudy (amistre64):

pi {S} 1 dx ; [0,ln3] becomes?

OpenStudy (anonymous):

pi (x)= pi (ln3)

OpenStudy (amistre64):

good :) 2pi - ln(3) pi is your answer then

OpenStudy (amistre64):

pi [ 2 - ln(3)] even lol

OpenStudy (anonymous):

and that is what the back of the book says...you are good lol

OpenStudy (amistre64):

toldja it was easy lol

OpenStudy (amistre64):

adding up areas of circles; thats all

OpenStudy (anonymous):

yea i just suck at it when my teacher tries to explain it..i may need more help in a few

OpenStudy (amistre64):

im sure well be around

OpenStudy (anonymous):

ok thank you :)

OpenStudy (anonymous):

u still there? i have a quick question?